40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 107 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  12237   Fri Jul 1 09:20:44 2016 SteveUpdateVACvent 78

  VENT OBJECTIVES:

1, Fix ETMX sus "jump issue"

2, First Contact clean the arms

3, Install new spare cold cathode and convectron gauges: InstruTech-Hornet

4, Install 50 mm apeture beam baffles

5, Check and clean optical quality viewport from inside

 

The following bullets were executed:

  • Closed shutter of PSL-IR and green shutters at the ends 
  • Checked jam nuts are protecting bellows 
  • Check crane functionality & cleanliness last week

  • Particle count must be under 10,000 counts / cf min for 0.5 micron 
  • Checked all metal window covers are on. 
  • Check 5 cylinders (24 cft size) of instrument grade air, called Alfa Gas 1 in stock.
  • Took pictures of medm screens: sus-sum, aligned oplevs pos, alignment values and vac configuration
  • Turned Oplev servos off
  • Closed V1 and VM1, opened VM2
  • Opened VV1 manual valve to Nitrogen cylinder and let the P1 to rise to 25 Torr
  • Switched over to Instrument Grade Air Cylinder and continoed the vent with 8-10 PSI reading at the pressure regulator

We are venting the 40m IFO

     

Quote:

Steve has ordered some teflon parts to take the place of the metal parts in his acetone-soaking jig. They should arrive tomorrow. 

So, we will be begin the venting process tomorrow. Doors to come off on Tuesday.

 

Attachment 1: vent78.png
vent78.png
Attachment 2: vent78at250T.png
vent78at250T.png
  12236   Fri Jul 1 01:52:54 2016 AakashUpdateGeneralSeismometer Enclosure Development | SURF 2016

I have transferred most of the temperature measurement stuff from the front area to seismometer at the end of Y-arm.  While arranging the components I have taken all care that they will not interfere with existing system. Also, I have temporarily taken a monitor from the front area to the area near same seismometer as I couldn't talk to Rpi via ssh. For next twelve hours, I am now recording temperature inside as well as outside the seismometer enclosure. Some temperature sensors are inside the enclosure while some are outside the seismometer enclosure.

 

  12235   Thu Jun 30 21:24:36 2016 ericqUpdateGeneralVent Prep

Steve has ordered some teflon parts to take the place of the metal parts in his acetone-soaking jig. They should arrive tomorrow. 

So, we will be begin the venting process tomorrow. Doors to come off on Tuesday.

  12234   Thu Jun 30 16:21:32 2016 gautamUpdateCOCSideband HOMs resonating in arms

[EricQ, gautam]

Last night, we set about trying to see if we could measure and verify the predictions of the simulations, and if there are indeed HOM sidebands co-resonating with the carrier. Koji pointed out that if we clip the transmitted beam from the arm incident on a PD, then the power of the higher order HG modes no longer integrate to 0 (i.e. the orthogonality is broken), and so if there are indeed some co-resonating modes, we should be able to see the beat between them on a spectrum analyzer. The procedure we followed was:

  1. Choose a suitable PD to measure the beat. We chose to use the Thorlabs PDA10CF because it has ~150MHz bandwidth, and also the responsivity is reasonable at 1064nm.
  2. We started our measurements at the Y-end. There was a sufficiently fast lens in the beam path between the transmon QPD and the high gain PD at the Y end, so we went ahead and simply switched out the high gain thorlabs PDA520 for the PDA10CF. To power the PDA10CF, we borrowed the power cable from the green REFL PD temporarily.
  3. We maximized the DC power of the photodiode signal using an oscilloscope. Then to introduce the above-mentioned clipping and orthogonality-breaking, we misaligned the beam on the PD until the DC power was ~2/3 the maximum value. 
  4. We then hooked up the PD output to the Agilent network analizyer (with a DC block).
  5. We measured the spectrum of the PD signal around 11.066MHz (with 100kHz span) and higher harmonics up to 55MHz and used a narrow bandwidth (100Hz) and long integration time (64 averages) to see if we could find any peaks. More details in the results section.
  6. Having satisfied ourselves with the Y-end measurements, we 
  • restored the power cable to the green beat PD
  • re-installed the thorlabs PDA520 
  • verified that both IR and green could be locked to the arm

We then repeated the above steps at the X-end (but here, an additional lens had to be installed to focus the IR beam onto the PDA10CF - there was, however, sufficient space on the table so we didn't need to remove the PDA520 for this measurement).


Results:

Y-end: DC power on the photodiode at optimal alignment ~ 200mV => spectra taken by deliberately misaligning the beam incident on the PD till the DC power was ~120mV (see remarks about these values).

RF sideband (Y-arm) Peak height (uV) Beat power (nW) RF sideband (X-arm) Peak height (uV) Beat Power (nW)
11 1.55 0.52 11 1.2 0.4
22 10.6 3.53 22 none seen N.A.
33 none seen N.A. 33 none seen N.A.
44 22.0 7.33 44 7 2.33
55 8.6 2.97 55 5 1.67

I converted the peak heights seen on the spectrum analyzer in volts to power by dividing by transimpedance (=5*10^3 V/A into a 50ohm load) * responsivity at 1064nm (~0.6A/W for PDA10CF).


Remarks:

  1. This effect flagged by the simulations seems to be real. Unfortunately I can't get a more quantitative picture because we can't quantify the mode-overlap between the carrier 00 mode and any higher order mode on the beat PD (as we know nothing about the profile of these modes), but the simulations did suggets that the 2nd order 22MHz and 4th order 44MHz HOMs are the ones closest to the carrier 00 resonance (see Attachments #2 and #3), which is kind of borne out by these results. 
  2. I disbelieve the conversions into power that I have done above, but have just put them in for now, because a DC power of 200mW at the Y-end suggests that there is >160uW of light transmitted from the arm, which is at least twice what we expect from a simple FP cavity calculation with the best-known parameters. If I've missed out something obvious in doing this conversion, please let me know! 
  3. For the Y-arm, the region around 55MHz had a peak (presumably from the sideband HOM beating with the carrier) but also a bunch of other weird sub-structures. I'm attaching a photo of the analyzer screen. Not sure what to make of this...
Attachment 1: image.jpeg
image.jpeg
Attachment 2: C1_HOMcurves_Y.pdf
C1_HOMcurves_Y.pdf
Attachment 3: C1_HOMcurves_X.pdf
C1_HOMcurves_X.pdf
  12233   Thu Jun 30 16:11:57 2016 ericqUpdateGeneralVent Prep

I have updated the vent prep checklist on the wiki. Gautam and I did the following things from it:

  • Center all oplevs, transmon QPDs 
    • ETMX oplev has not been centered, since it's moving around so much, and we're going to immediately move the suspension anyways.
  • Align the arm cavities for IR and align the green lasers to the arms. 
    • AUX X Green was aligned while the X arm was well aligned. Soon thereafter, ETMX wandered away, but the green will remain a good reference
  • Update the SUS Dritmon values 
  • Reconcile all SDF differences 
  • Reduce input power to no more than 100mW by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.) 

  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
    • I don't think we've vented since the most recent slew of changes to the IMC servo, so its not surprising that the current low power scripts don't work. I'm working on locking the IMC, but this does not prevent us from initiating the vent tomorrow.

 

The following bullets have not yet been executed:

  • Close shutter of PSL-IR and green shutters at the ends 
  • Make sure the jam nuts are protecting bellows 
  • Check crane functionality & cleanliness 

  • Turn off HV into vacuum: OMC is not wired this time 
  • Particle count must be under 10,000 counts / cf min for 0.5 micron 
  • Check all metal window covers are on. 
  • Check 5 cylinders (24 cft size) of instrument grade air, called Alfa Gas 1 in stock.
  12232   Thu Jun 30 14:31:02 2016 ChemistryUpdateSUSNO

  12231   Thu Jun 30 10:03:16 2016 SteveUpdateSUS how to remove epoxy from optics

Proposed Acetone soak dish for SOS epoxy softening.

It has good acces through 5" top ID. The set up is stable and teflon lined.

Materials: glass jar with SS cover, teflon bricks, 0.008" teflon wrapped "high density Drever bricks" and aluminum

Drever brick: I beleive it is a Tungsten alloy. We used it as vac-bat savor at the coffe can. It has high density, heavy and hard, it was never identified.

I will soak one brick  to see if it has any reaction ability with acetone.

 

NO means that only Glass and Teflon can be used for this fixture in  Acetone. We can not take a chance on the coating!

I guess the small surface area Aluminum dumbbell, guide rod and-or wire standoff, magnet and epoxy does not degrade the acetone such way that it effects our coating.

Not ot mention, that only the very edge of the coating would in this solution.

              

 

Attachment 1: SOSsoakDish.jpg
SOSsoakDish.jpg
Attachment 2: SOSsoakDish_tv.jpg
SOSsoakDish_tv.jpg
  12230   Wed Jun 29 16:05:47 2016 ericqUpdateSUSETMX suspension readiness

I've gone through the SOS suspension document (E970037) and some old elogs to get an idea of all the accesories we need for the process of suspending, aside from the tower itself, which Steve has already put together. Gautam and I have laid our eyes upon most of the critical pieces. Some other objects are unknown, and perhaps not strictly neccesary. 

Confirmed to exist:

  • Guide rod gluing fixture
  • 3 axis micrometer + microscope doohicky
  • PZT buzzer (for fine standoff positioning)
  • Winch fixture (the plate maybe doesn't fit perfectly on the top of the SOS tower, but one bolt will probably work)
  • Spare guide rods and ruby standoffs
  • Magnet+dumbell gluing fixture (in case we need to glue a new magnet+dumbell pair)
  • Magnet to optic gluing fixture (including "pickle pickers"/grippers)
  • 0.017" suspension wire
  • HeNe laser + QPD + beam height target (confirmed working on scope)

In addition, I am told that we have a long ribbon cable that can run from the X end to the clean room to enable OSEM damping control while we do the pitch alignment.

Things mentioned in the procedure I have not found:

  • Something called an "SOS set screw tool" (D961412)
    • Not mentioned specifically what this is for
  • Edmund scientific pocket measuring microscope + bushing
    • Seems to be intended for centering OSEM holding plate on the magnet position. I found no mention of this use in 40m elogs. It can likely be done by eye
  • SOS cleaning bracket (D970181)
    • This is likely for the old liquinox cleaning procedure. We'll be using first contact.

Some other tasks and their status:

  • Check cleanroom table levelling:
    • Done via 12" spirit level, looks good
  • Glue removal strategy / fixture
    • Steve is working on this
  • Cut and apply viton tips to earthquake stops
    • Either gautam will do this, or we will use the current suspension's stops
  12229   Wed Jun 29 16:02:29 2016 SteveUpdateSUS ETMX beam baffle

In the attached photo from 2012, one can see the installed black glass baffle. According to the drawings (LIGO-DNNNXXX) this one has a clear aperture of 40 mm.

In (someplace ?) we have clean baffles with a 50 mm aperture which can be installed during this vent. In order to be more conservative, let us choose to swap these out for all 4 test masses during the upcoming vent using the green laser as an alignment guide, as Koji described at today's lunch meeting.

They are located at the top of E1 drawer cabinet

Attachment 1: IMG_1608.JPG
IMG_1608.JPG
  12228   Wed Jun 29 15:34:00 2016 SteveUpdatesafetySURF 2016 safety

Praful Vasiceddy received 40m specific basic safety training.

Quote:

Aakash Patil received 40m specific basic safety training.

Quote:
Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

Varun has received 40m specific basic safety training today.

 

 

  12227   Wed Jun 29 14:32:22 2016 SteveUpdateSUS how to remove epoxy from optics

 

Hi Steve -

I found the doc I was looking for:

https://dcc.ligo.org/LIGO-E1200821

Specifically, you might find guidance in Section 5 and the pictures at the end of the doc.  This should work for Vacseal as well.

Good luck - it will take some time (hours to day or 2)...
I'll be interested to know how it goes.

GariLynn helped us develop this procedure so you could also ask her to cast an eye over the setup if you are worried.

-Betsy

ps: there is no existing fixture to hold SOS optic while soaking it
 

  12226   Wed Jun 29 08:20:04 2016 SteveUpdateSUS First Contact shelf life
We'll follow LIGO policy:
Our policy is to use first contact within 1 year of purchase for use in the interferometers.  For inspection use I am comfortable with out-of-date use.

GaryLinn offered their indate First Contact for use.

 

  12225   Wed Jun 29 00:09:36 2016 AakashUpdateGeneralThings from past | SURF 2016

I have taken out the heaters and temperature sensors from the enclosure which was made by Megan last summer. Soon I will test and configure those heaters.

  12224   Tue Jun 28 22:54:43 2016 AakashUpdateGeneralSeismometer Enclosure Development | SURF 2016

The existing enclosure for seismometer at LIGO 40m lab is a cylindrical stainless steel can placed upside down over the seismometer. It has more empty space between the seismometer and the internal surface of enclosure which is not desirable(I'll quantitatively elaborate this statement once my temperature measuring setup is ready).

 

Stainless steel has a thermal conductivity in the range of 16.3 to 16.7 W/m/K and magnetic permeability 1.260e-6 H/m.Assuming an ambient temperature 298K, and the temperature inside the enclosure as 295K, as well as substituting all the values for dimesions and material properties of existing enclosure,
k=16.4 W/mK, μ=1.260e-6 H/m, L=2ft=0.6096m, b=r2 =0.5ft=0.1524m, thickness=5mm, a=r1 =0.1474m.
So by using the textbook relations(I have mentioned them in my report), the value of attenuation coefficient is 5.953584e-05 and the value of rate of heat transfer= 5.64913 kW. The attenuation coefficient value is quite better for steel but proper care needs to be taken to avoid heat transfer. For studying the variation of rate of heat transfer and attenuation with the thickness of enclosure material, I have plotted the following attached graphs for different materials which include hardened stainless steel, aluminium, pure iron and nanoperm-muMetal.

 

 

About Data Acquisation

I have already invested a lot of time to configure and use acromag busworks card over ethernet. So now I have made an arrangement to measure temperature by AD592CNZ temperature transducer IC. I would be using raspberry pi for acquiring data untill I figure out a way to use acromag busworks card for the same. This setup of acquiring logging temperature using raspberry pi is mostly ready except the calibration part.

  12223   Tue Jun 28 20:43:23 2016 KojiSummaryCOCFirst Contact cleaning practice

Made a dry run of the in-situ cleaning for a 3inch optic.

Attachment 1: The Al dummy mass is clamped in the suspension cage.
Attachment 2: The front surface was painted. The nominal brush with the FC bottle was used.
Attachment 3: Zoom in of the front surface.
Attachment 4: The back surface was painted.
Attachment 5: The back surface was peeled.
Attachment 6: The front surface was peeled too.
Attachment 7: The peeled layers.

Findings:

1. To paint a thick layer (particlarly on the rim) is the key to peel it nicely.

2. It was helpful for easier peeling to have mutiple peek tabs. Two tabs were sufficient for ~1" circle.

3. The nominal brush with the bottle was OK although one has to apply the liquid many times to cover such a large area. A larger brush may cause dripping.

4. The nominal brush was sufficiently long once the OSEMs are removed. In any case it is better to remove the OSEMs.

Attachment 1: IMG_20160628_170335196.jpg
IMG_20160628_170335196.jpg
Attachment 2: IMG_20160628_171547769.jpg
IMG_20160628_171547769.jpg
Attachment 3: IMG_20160628_171607802.jpg
IMG_20160628_171607802.jpg
Attachment 4: IMG_20160628_172328190.jpg
IMG_20160628_172328190.jpg
Attachment 5: IMG_20160628_174541960.jpg
IMG_20160628_174541960.jpg
Attachment 6: IMG_20160628_174556004.jpg
IMG_20160628_174556004.jpg
Attachment 7: IMG_20160628_174617198.jpg
IMG_20160628_174617198.jpg
  12222   Tue Jun 28 17:11:27 2016 PrafulUpdateGeneralEM172 Microphones

Found 60 EM172 microphones. Previous elog with details: 7777.

  12221   Tue Jun 28 16:10:49 2016 PrafulUpdateGeneralBluebird Microphones

Found 1 out of 2 bluebird microphones in the 40m.

  12220   Tue Jun 28 16:09:41 2016 PrafulUpdateGeneral40m Summary Pages

Set up gwsumm on optimus and generated summary pages from both L1 and C1 data. Still a few manual steps need to be taken during generation, not fully automated due to some network/username issues. nds2 now working from optimus after restarting nds2 server.

  12219   Tue Jun 28 16:06:09 2016 gautamUpdateCOCRC folding mirrors - further checks

Having investigated the mode-overlap as a function of RoC of the PRC and SRC folding mirrors, I've now been looking into possible stability issues, with the help of some code that EricQ wrote some time back for a similar investigation, but using Finesse to calculate the round trip Gouy phase and other relevant parameters for our current IFO configuration. 

To do so, I've been using:

  1. Most up to date arm length measurements of 37.81m for the Y arm and 37.79m for the X arm
  2. RoCs of all the mirrors from the phase map summary page
  3. Loss numbers from our November investigations

As a first check, I used flat folding mirrors to see what the HOM coupling structure into the IFO is like (the idea being then to track the positions of HOM resonances in terms of CARM offset as I sweep the RoC of the folding mirror). 

However, just working with the flat folding mirror configuration suggests that there are order 2 22MHz and order 4 44MHz HOM resonances that are really close to the carrier resonance (see attached plots). This seems to be originating from the fact that the Y-arm length is 37.81m (while the "ideal" length is 37.795m), and also the fact that the ETM RoCs are ~3m larger than the design specification of 57m. Interestingly, this problem isn't completely mitigated if we use the ideal arm lengths, although the order 2 resonances do move further away from the carrier resonance, but are still around a CARM offset of +/- 2nm. If we use the design RoC for the ETMs of 57m, then the HOM resonances move completely off the scale of these plots... 

Attachment 1: C1_HOMcurves_Y.pdf
C1_HOMcurves_Y.pdf
Attachment 2: C1_HOMcurves_DR.pdf
C1_HOMcurves_DR.pdf
  12218   Tue Jun 28 09:14:27 2016 SteveUpdateSUS ETMX optics preparation to be rehanged

Given the effect on the contrast defect, the consensus at the meeting Wednesday 6-22  was that we should continue to use the existing ETMX optic. 

 
Does this imply that we can simply:
  • Vent
  • Remove the current ETMX suspension, and optic from the suspension
  • Replace the standoffs (and guide rod if necessary), leaving the current magnets in place
  • Re-suspend on the current, existing suspension, with new wire clamps
That is, if we are keeping the current optic, do we gain anything by replacing the whole suspension? Is there any reason to remove and replace magnets? (The wire clamps do need to be replaced, as a groove forms when they are installed.)
 
This could allow us to vent and execute this whole thing sooner. 
 
-Eric Q.
...............................
Yes, I think this is a reasonable approach. Besides the wire standoffs and wireclamps, the only other problem might by a worn groove in the top piece of the suspension (not the clamp piece, but the piece that the clamp holds the wire up against).
 
How about use the current ETM, but put it into the new suspension body with new wire standoffs?
 
Quote:

D Location

Number on

Drawing

 

Component

Name

 

 Baked Clean

 Pieces Needed

 

Pieces

In Stock

(on tower )

Notes
         
31,20,19,13 viton tips   not cut yet baked material in stock
30 6-32x0.75" stops 4 (4)  
29 steel music wire 0.017" not baked on roll  needs good wipes
28  1/4 "washer 4 (4)  
27 lock washer 4 50 install
26 Ag plated 1/4-20x1.25 4 (4)  
25 Ag plated1/4-25x.75 & not plated 20 (20)  
24 SS 4-40x.5 2 (2)  
23 SS 4-40x.38 4 (4)  
22 spring plunger 4+2 (4+2)  
         
         
18 magnets, 1.9mm od, length 3.2 mm 5 ~30 not coated, rusty buy Ni coated ones for future use
17 guide rod, 0.635 mm od, 3.3 mm 3 6 Al
16 wire standoff, 1 mm od, 4.8 mm 2 2 Al and ruby (ruby groove not centered)
15 short OSEMs 5 6  
14 spare ETMX in 40m wiki 1 1  confirmed in cabinet
         
12 dumbbell standoff 5 6  
11 Al stiffening plate 1 (1)  
10 wire clamp B in sus block 2 (2)  
9 wire clamp A 1 (1)  
8-7 lower clamp for lifting optics 2 (2)  
6 upper clamp to hold down optic 1 (1)  
5-4 left-right side of tower  1 ea (1ea)  
3 tower base 1 (1)  
2 sus block 1 (1)  
1 lower and upper OSEM holders 1ea (1ea)  
         
48 sandind fixture for magnet&dumbbell      
45 magnet-dumbbell assemblly fixture 1 1  
43 guide rod gluing fixture     shipped from LOH .... D000335...needs cleaning
         
  First contact   3-15-2013 purchase check expiration date OK
  FCPEEK peeler ring disk for TM cleaning 10 front,10 back side  have sheets only 32&19mm ID punches ordered
  GordonBrush custom for LIGO optical cleaning ~5 1 (3/8wide nylonSS) more from Calum available
  EP30-2 epoxy have  have expiration date 9-24-2016

NOT finished, last edited 6-28

 

Attachment 1: nylonFCss.jpg
nylonFCss.jpg
  12217   Mon Jun 27 15:47:17 2016 SteveUpdateGeneralQPR clean room gloves

I got some QPR Nitrile gloves. They are LIGO approved.White nitrile gloves are naturally anti-static- 109 ohms

Their touch not as good as laytex  gloves but try to use them.

 

  12216   Mon Jun 27 15:26:03 2016 SteveOmnistructureALARMfire alarm test

The fire alarm came on around 15:05  for about 2-3 minutes. We all  left the lab and counted heads.  I called Paul Mackel x2646 (cell 626/ 890- 3259) at Fire Protection Services. He said that this alarm test was planned and we should of got an email notice. Perhaps I missed that notes.

Quote:

Fire alarm went off several minutes ago. Talked to security and they said there was no fire. It beeped twice again just now. No one has been working on the IFO today.

 

Attachment 1: fireAlarmTest.png
fireAlarmTest.png
  12215   Mon Jun 27 15:12:09 2016 varunUpdateCDSDAFI update: stereo output

Using an RC to BNC connector from the inner drawer, I have added a second output cable going from the output Fibox in the control room to the audio mixer.

Quote:

I have updated the DAFI with the following changes:

1) Separated both the channels of stereo output completely, as well as in the GUI.

2) Added text monitors for the inputs and outputs.

The stereo output is now ready except for a cable going from the second channel of the output fibox to the audio mixer.

Attached is the main DAF_OVERVIEW screen and its link button from the LSC screen labelled "DAFI"

Quote:

I wish to have stereo audio output for the DAF module. Hence, there needs to be a second output from the DAF. I added this second output to the model. Following are the details:

FiBox: It consists of two analog inputs which are digitized and multiplexed and transmitted optically. (only 1 fiber is needed due to multiplexing). Attachment 1 shows the fibox with its 2 analog inputs (one of which, is connected), and 1 fiber output. The output of the DAF goes to the FiBox. Until today, the Fibox recieved only 1 analog input. This analog signal comes from the DAC-8 (count starting from 0), which is located at "CH 1 OUT" SMA output in the "MONITORS" bin on the racks (attachment 2).

I have added another output channel to the DAF model both in software and in hardware. The DAF now also uses DAC-9 analog output which goes to the second analog input of the FiBox. The DAC-9 output is located at "CH 2 OUT" SMA output in the "MONITORS" bin on the racks (attachment 4).

After making the changes, the Fibox is shown in attacment 3.

Testing: The LSC input on passing through the DAF block is given through two different DAC outputs, to the same Fibox channel (one after the other), and the output is heard. More concrete testing will be done tomorrow. It will be as follows:

1) Currently, I need to search for a suitable cable that would connect the second channel of the output fibox to the audio mixer. After doing this, end to end testing of both channels will be done.

2) I could not access the AWG, probably because the DAQ was offline today afternoon. Using a signal from the AWG will give a more concrete testing of the stereo output.

3) After this, I will separate the two channels of the stereo completely (currectly they are seperated only at the DAF output stage)

4) I also will edit the medm gui appropriately.

 

Quote:

I have added Enable buttons for each of the DSP blocks, and labels for the matrix elements. The input matrix takes inputs from each of the 4 channels: ADC1, ADC2, LSC and EXC, and routes them to the audio processing blocks (attachment 2). The output matrix (attachment 3) takes the outputs of the various DSP blocks and routes them to the output and then to the speakers. 

 

 

 

Attachment 1: IMG_20160627_151753247.jpg
IMG_20160627_151753247.jpg
  12214   Sun Jun 26 15:27:28 2016 ranaFrogsIOOPMC /MC lopced

Found PMC unlocked for many hours so I relocked it. IMC relocked by itself, but the input switch seems to be flickering to fast. Also the Keep Alive bit is not flashing. no

  12213   Thu Jun 23 17:24:32 2016 ericqUpdateGeneralTweaks

I spent some time this afternoon reviving some of my CESAR/ESCOBAR shenanigans on the Y arm. I found it neccesary to adjust a few things.

  • PMC realigned 
  • ETMY oplev centered
  • Y End green realigned
  • PSL/Y Green beat realigned

Afterwards, ALSY noise levels were good. 

  12212   Wed Jun 22 14:03:42 2016 VarunUpdateElectronicsAnti-Aliasing Filter circuit schematic

I found an anti-aliasing circuit on the 40m wiki. It consists of A differential LPF made using THS4131 low noise differential op-amp (one of the main applications of which is preprocessing before the ADC), and a notch. I modified it to arrange for the desired bandwidth (about 8 kHz) and notch after the Nyquist frequency at 36 kHz. I simulated it to get the attached results:

Attachment 1: It shows the input PSD (same as the one posted in the previous elog), the filter transfer function, and The resulting output.

Attachment 2: The circuit schematic. The initial part using THS4131 is a differential LPF and the subsequent RC network is the notch.

Attachment 3: This shows the ratio of the aliased downconverted signal to the the in-band signal, representative of the contamination in each bin. Here too, the aliased signals are negligible as compared to the low frequencies but they are not negligible as compared to the higher frequencies (above 10 kHz) into which they would get downconverted due to sampling. However, here, the attenuation at 8kHz is less than 6 dB while in the previous circuit, it was about 12 dB. One problem with this circuit is at about 6kHz, there is aliased signal from the 65k to 98kHz band, but this can be taken care of by adding an LPF later.

Quote:

Summary: The aim is to design an analog anti-aliasing (AA) filter placed before the ADC, whose function is to filter out components of the input spectrum that have frequencies higher than the Nyquist frequency. This needs to be done so that there is no contamination of aliased downconverted high-frequency signals into the ADC output. I have put down and simulated a circuit to do this, based on the spectra of a few interferometer signals that eric Provided. Attachment 1 shows such an input PSD, treated with whitening filter, before the AA. The sampling rate is 65536 Hz and hence the Nyquist freq. is 32768 Hz.

Motivation: Attachments 2 and 3 show the plot of required attenuation for various frequencies above the Nyquist. We can see a peak at 36 kHz, which will alias to about 29kHz. It will require about 70 dB attenuation here. This indicates that use of a notch filter combined with a low pass filter can be used.

Details of Schematic: Attachment 4 shows the schematic of a Boctor low pass notch filter, cascaded by a 2nd order LPF. The stopband frequency of the boctor filter can be tuned to around 36 kHz. Its main advantage for the boctor is better insensitivity to component value tolerances, use of a single op amp, and relatively independent tuning of parameters.  The various component values are calculated from here. The transfer functions for the circuit shown in attachment 4 were simulated using TINA - a spice based simulation software. The transfer function is shown in attachment 5.

A few more calculations: Attachment 6 shows the output psd after the signal has been treated with AA. Attachments 7 and 8 show the ratio of aliased downconverted signal and the unaliased signal of the output. Here, we can see that above about 13 kHz, the ratios go above -40dB, which is apparently undesirable. However, we also see from the transfer function of the filter that the gain falls to less than -20dB after about this frequency, and the aliased signals are atleast 20 dB lower than this, atleast upto about 29 kHz in attachment 7 and about 25 kHz in attachment 8. This means that the aliased signals are negligible as compared to the low frequencies even if they are not negligible as compared to the higher frequencies (above 13 kHz) into which they would get downconverted due to sampling. But these higher frequencies (above 13 kHz) themselves are small.

The filter overall, is 4th order. Considering this and the above discussion, I need to decide what changes to make in the existing schematic. For now, I could discuss with eric to finalize the opamp and start building the pcb board design.

 

Attachment 1: io.pdf
io.pdf
Attachment 2: AA.JPG
AA.JPG
Attachment 3: ratios_v2.pdf
ratios_v2.pdf
  12211   Wed Jun 22 10:15:45 2016 varunUpdateCDSDAFI update: stereo output

I have updated the DAFI with the following changes:

1) Separated both the channels of stereo output completely, as well as in the GUI.

2) Added text monitors for the inputs and outputs.

The stereo output is now ready except for a cable going from the second channel of the output fibox to the audio mixer.

Attached is the main DAF_OVERVIEW screen and its link button from the LSC screen labelled "DAFI"

Quote:

I wish to have stereo audio output for the DAF module. Hence, there needs to be a second output from the DAF. I added this second output to the model. Following are the details:

FiBox: It consists of two analog inputs which are digitized and multiplexed and transmitted optically. (only 1 fiber is needed due to multiplexing). Attachment 1 shows the fibox with its 2 analog inputs (one of which, is connected), and 1 fiber output. The output of the DAF goes to the FiBox. Until today, the Fibox recieved only 1 analog input. This analog signal comes from the DAC-8 (count starting from 0), which is located at "CH 1 OUT" SMA output in the "MONITORS" bin on the racks (attachment 2).

I have added another output channel to the DAF model both in software and in hardware. The DAF now also uses DAC-9 analog output which goes to the second analog input of the FiBox. The DAC-9 output is located at "CH 2 OUT" SMA output in the "MONITORS" bin on the racks (attachment 4).

After making the changes, the Fibox is shown in attacment 3.

Testing: The LSC input on passing through the DAF block is given through two different DAC outputs, to the same Fibox channel (one after the other), and the output is heard. More concrete testing will be done tomorrow. It will be as follows:

1) Currently, I need to search for a suitable cable that would connect the second channel of the output fibox to the audio mixer. After doing this, end to end testing of both channels will be done.

2) I could not access the AWG, probably because the DAQ was offline today afternoon. Using a signal from the AWG will give a more concrete testing of the stereo output.

3) After this, I will separate the two channels of the stereo completely (currectly they are seperated only at the DAF output stage)

4) I also will edit the medm gui appropriately.

 

Quote:

I have added Enable buttons for each of the DSP blocks, and labels for the matrix elements. The input matrix takes inputs from each of the 4 channels: ADC1, ADC2, LSC and EXC, and routes them to the audio processing blocks (attachment 2). The output matrix (attachment 3) takes the outputs of the various DSP blocks and routes them to the output and then to the speakers. 

 

 

Attachment 1: C1DAF_OVERVIEW.png
C1DAF_OVERVIEW.png
Attachment 2: DAF_link_from_LSC.png
DAF_link_from_LSC.png
  12210   Wed Jun 22 08:40:42 2016 ranaUpdateLSCY arm @ 30kHz UGF w/POY, AO

Below 100 Hz, I suppose this means that the X arm is now limited by the quadrature sum of the X and Y arm seismic noise.

  12209   Tue Jun 21 14:12:10 2016 SteveUpdatePEM vacuum envelope wiped

I wiped down the cranes with wet towel and Mario our janitor did the chamber tops with the tubes.

The optical tabels were not touched.

The outside temp peaked at 44 C yesterday

 

Attachment 1: 60dPEM.png
60dPEM.png
  12208   Tue Jun 21 11:49:29 2016 ericqFrogsCDSmedm command not working

The workstations' .bashrc is a symbolic link to /users/controls/.bashrc

In it, someone commented out the critical line:

#source /ligo/cdscfg/workstationrc.sh

I uncommented it. medm (and all of the other things like cdsutils) work again.

I blame jamie.

  12207   Tue Jun 21 11:26:42 2016 varunFrogsCDSmedm command not working

"medm: command not found" error when run through command line both in pianosa and rossa in both editing and execution modes. It however gets executed and edited through the sitemap button. Don't know the source of the problem. Gautam did check the .bashrc file. aliases for SITEMAP and m40m are intact in the .bashrc file.

  12205   Tue Jun 21 04:01:09 2016 ericqUpdateLSCY arm @ 30kHz UGF w/POY, AO

With the newly amplified POY signal, locking the mode cleaner to the Y arm at ~30kHz bandwidth was quite straightforward. The offset jumps still happen, and are visible in POY11_I_ERR, but are never big enough to cause much power degradation in TRY (except when turning on CM board boosts, but its still not enough to lose lock). The script which accomplishes this is at scripts/YARM, and is in the svn. The MC2/AO crossover is at about 150Hz with 40deg margin.

For now, I'm using IN1 of the CM board, because I haven't removed the op27s that I put into IN2's gain stages. I believe the slew rate limitations of these prevent them from working completely during the offset jumps. I'll put AD829s back soon. 

At first, I had ITMX misalgined to use AS55 as an out of loop sensor, then I aligned and locked the X arm on POX to compare.

Weirdly enough, locking the mode cleaner to the Y arm with 30kHz UGF and two boosts on make no real visible difference in the X arm control signal. This is strange, as the whole point of this affair was to remove the presumably large influence of frequency noise on the X arm signals... Maybe this is injecting too much POY sensor noise?

 

Attachment 1: inloop_outloop.pdf
inloop_outloop.pdf
Attachment 2: xarm.pdf
xarm.pdf
  12204   Mon Jun 20 18:07:15 2016 gautamUpdateCOCContrast as a function of RoC of ETMX
Quote:

So, it seems that changing the ETMX for one of the spares will change the contrast defect from ~0.1% to 0.9%. True? Seems like that might be a big deal.

That is what the simulation suggests... I repeated the simulation for a PRFPMI configuration (i.e. no SRM, everything else  as per the most up to date 40m numbers), and the conclusion is roughly the same - the contrast defect degrades from ~0.1% to ~1.4%... So I would say this is significant. I also attempted to see what the contribution of the asymmetry in loss in the arms is, by running over the simulation with the current loss numbers of 230ppm for Yarm and 484ppm for the X arm, split equally between the ITMs and ETMs for both cases, and then again with lossless arms - see attachment #1. While this is a factor, this plot seems to suggest that the RoC mismatch effect dominates the contrast defect...

Attachment 1: contrastDefectComparison.pdf
contrastDefectComparison.pdf
  12203   Mon Jun 20 16:33:09 2016 VarunUpdateElectronicsAnti-Aliasing Filter circuit schematic

Summary: The aim is to design an analog anti-aliasing (AA) filter placed before the ADC, whose function is to filter out components of the input spectrum that have frequencies higher than the Nyquist frequency. This needs to be done so that there is no contamination of aliased downconverted high-frequency signals into the ADC output. I have put down and simulated a circuit to do this, based on the spectra of a few interferometer signals that eric Provided. Attachment 1 shows such an input PSD, treated with whitening filter, before the AA. The sampling rate is 65536 Hz and hence the Nyquist freq. is 32768 Hz.

Motivation: Attachments 2 and 3 show the plot of required attenuation for various frequencies above the Nyquist. We can see a peak at 36 kHz, which will alias to about 29kHz. It will require about 70 dB attenuation here. This indicates that use of a notch filter combined with a low pass filter can be used.

Details of Schematic: Attachment 4 shows the schematic of a Boctor low pass notch filter, cascaded by a 2nd order LPF. The stopband frequency of the boctor filter can be tuned to around 36 kHz. Its main advantage for the boctor is better insensitivity to component value tolerances, use of a single op amp, and relatively independent tuning of parameters.  The various component values are calculated from here. The transfer functions for the circuit shown in attachment 4 were simulated using TINA - a spice based simulation software. The transfer function is shown in attachment 5.

A few more calculations: Attachment 6 shows the output psd after the signal has been treated with AA. Attachments 7 and 8 show the ratio of aliased downconverted signal and the unaliased signal of the output. Here, we can see that above about 13 kHz, the ratios go above -40dB, which is apparently undesirable. However, we also see from the transfer function of the filter that the gain falls to less than -20dB after about this frequency, and the aliased signals are atleast 20 dB lower than this, atleast upto about 29 kHz in attachment 7 and about 25 kHz in attachment 8. This means that the aliased signals are negligible as compared to the low frequencies even if they are not negligible as compared to the higher frequencies (above 13 kHz) into which they would get downconverted due to sampling. But these higher frequencies (above 13 kHz) themselves are small.

The filter overall, is 4th order. Considering this and the above discussion, I need to decide what changes to make in the existing schematic. For now, I could discuss with eric to finalize the opamp and start building the pcb board design.

Attachment 1: in.pdf
in.pdf
Attachment 2: 32to65att.pdf
32to65att.pdf
Attachment 3: 65to98att.pdf
65to98att.pdf
Attachment 4: lpf_notch.JPG
lpf_notch.JPG
Attachment 5: lpf_notch.pdf
lpf_notch.pdf
Attachment 6: out.pdf
out.pdf
Attachment 7: out_ratio1.pdf
out_ratio1.pdf
Attachment 8: out_ratio2.pdf
out_ratio2.pdf
  12202   Mon Jun 20 14:03:04 2016 jamieConfigurationCDSEndRun GPS receiver upgraded, fixed

I just upgraded the EndRun Technologies Tempus LX GPS receiver timing unit, and it seems to have fixed all the problems.  cool

Thanks to Steve for getting the info from EndRun.  There was indeed a bug in the firmware that was fixed with a firmware upgrade.

I upgraded both the system firmware and the firmware of the GPS subsystem:

Tempus LX GPS(root@Tempus:~)-> gntpversion 
Tempus LX GPS 6010-0044-000 v 5.70 - Wed Oct 1 04:28:34 UTC 2014
Tempus LX GPS(root@Tempus:~)-> gpsversion 
F/W 5.10 FPGA 0416
Tempus LX GPS(root@Tempus:~)->

After reboot the system is fully functional, displaying the correct time, and outputting the correct IRIG-B data, as confirmed by the VME timing unit.

I added a wiki page for the unit: https://wiki-40m.ligo.caltech.edu/NTP

 

Steve added this picture

Attachment 1: GPSreceiverWorking.jpg
GPSreceiverWorking.jpg
  12201   Mon Jun 20 11:19:41 2016 jamieConfigurationCDSGPS receiver not resetting properly
Quote:

I called https://www.endruntechnologies.com/pdf/USM3014-0000-000.pdf  and they said it's very likely just needs a software update. They will email Jamie the details.

I got the email from them.  There was apparently a bug that manifested on February 14 2016.  I'll try to software update today.

http://endruntechnologies.com/pdf/FSB160218.pdf
http://endruntechnologies.com/upgradetemplx.htm

  12200   Mon Jun 20 11:11:20 2016 SteveConfigurationCDSGPS receiver not resetting properly

I called https://www.endruntechnologies.com/pdf/USM3014-0000-000.pdf  and they said it's very likely just needs a software update. They will email Jamie the details.

Quote:

The GPS receiver (EndRun Technologies box in 1Y5? (rack closest to door)) seems to not coming back up properly after the reboot.  The front pannel says that it's "LKD", but the "sync" LED is flashing instead of solid, and the time of year displayed on the front panel is showing day 6.  The fb1 symmetricom driver and VME timing module are still both seeing day 299, though.  So something may definitely be screwy with the GPS receiver.

 

  12199   Mon Jun 20 09:36:10 2016 SteveUpdateSUS EQ 3.1 m Big Bear Lake, CA

No suspention lost damping.

 

Attachment 1: eq3.1BigBearLakeCA.png
eq3.1BigBearLakeCA.png
  12198   Mon Jun 20 08:26:56 2016 SteveUpdatePSLAOM pictures

Good job Johannes and Subham.

 

Attachment 1: AOMinplace.jpg
AOMinplace.jpg
Attachment 2: inplaceAOM.jpg
inplaceAOM.jpg
Attachment 3: AOMin.png
AOMin.png
  12197   Mon Jun 20 01:38:04 2016 ranaUpdateCOCContrast as a function of RoC of ETMX

So, it seems that changing the ETMX for one of the spares will change the contrast defect from ~0.1% to 0.9%. True? Seems like that might be a big deal.

  12196   Fri Jun 17 22:36:11 2016 JohannesUpdatePSLAOM installation

Subham and I have placed the AOM back into the setup right in front of the PMC.

Steps undertaken:

  1. The HEPA filters were turned off for some reason. They were turned back on, running at 100% while the enclosure was open.
  2. Before the installation, after initial realignment, the PMC TRANSPD read out 748 mV.
  3. The laser injection current was dialed down to 0.8 A (just above the threshold, judging by PMC cameras.
  4. AOM was attached to the new mount while staying connected to its driver. Put in place, a clamp prevents the cable from moving anywhere near the main beam.
  5. Aligned AOM to beam, centering the beam (by eye) on front and back apertures.
  6. We then applied an offset to the AOM driver input, eventually increasing it to 0.5 V. A secondary beam became clearly visible below the primary beam.
  7. In order to place the razor blade dump (stemming from a box, labeled "cleaned for atm use") before the PMC, where the beam separation was about 3 mm, to make sure we can hit the edged area, we had to place the dump at an angle, facing up.
  8. Keeping the 0.5V offset on the driver input, with the lights off, we increased the laser diode current in steps of ~200 mA to its original value of 2.1A, while checking for any IR light scattered from the beam dump. Not a trace.
  9. At original current setting, we realigned the beam into the PMC, and obtained 743 mV on the TRANSPD in the locked state.
  10. Closed off PSL table, dialed HEPAs down to 50%

              

 

Attachment 1: aom_new_mount.jpg
aom_new_mount.jpg
  12195   Fri Jun 17 15:22:31 2016 gautamUpdateVACN2 supply line restored after retiling
Quote:

The drill room floor will be retiled Thursday, June 16. Temporary nitrogen line set up will allow emptying the hole area.

 

Ifo room entry will be through control room.

 

The retiling work has finished, Steve and I restored the N2 supply configuration to its normal state. The sequence of steps followed was:

  1. Went to the X end and closed the following valves, roughly in this order: VAEE, VAEV, VABS, VABSCI, VASV, VASE, V4, V1.
  2. Checked the RPM on the various turbo pump controllers to make sure they were in their nominal states
  3. Disconnect the electrical connections to V1, V4, V5 and VA6 - just to make sure some spurious signal doesn't unintentionally open any of these valves while we are mucking around with the N2 supply
  4. Close the valves on the N2 cylinders in the drill room. Disconnect the temporary nitrogen line (at this point, the N2 pressure to the IFO valves goes down from ~7-PSI to 0), reconnect the old supply chain, taking care that we aren't unintentionally loosening any of the Swagelock connections while unscrewing stuff
  5. Replaced one of the N2 cylinders that was running low.
  6. Reopen the cylinders, restore N2 pressure to IFO valves to ~70PSI.
  7. Do steps 1-3 in reverse: i.e. reconnect power to all valves, open them in the reverse order we closed them while monitoring the state of the various turbo pumps. 
  8. Acknowledged the error message on the C0VAC medm screen

Note: the valve isolating the RGA automatically shutoff during this work, possibly because it detected a pressure above its threshold - after checking the appropriate pressure gauges, we reopened this valve as well. 

The attached screenshot suggests that everything went as planned and that the vacuum system is back to normal...

 

 

Attachment 1: c0vac_06172016.png
c0vac_06172016.png
  12194   Thu Jun 16 23:02:57 2016 gautamUpdateCOCContrast as a function of RoC of ETMX
Quote:

That sounds weird. frownIf the ETMY RoC is 60 m, why would you use 57.6 m in the simulation? According to the phase map web page, it really is 60.2 m.

This was an oversight on my part. I've updated the .kat file to have all the optics have the RoC as per the phase map page. I then re-did the tracing of the Y arm cavity mode to determine the appropriate beam parameters at the laser in the simulation, and repeated the sweep of RoC of ETMX while holding RoC of ETMY fixed at 60.2m. The revised contrast defect plot is attached (this time it is the contrast defect, and not the contrast, but since I was running the simulation again I thought I may as well change the plot). 

As per this plot, if the ETMX RoC is ~54.8m (the closer of the two spares to 60.2m), the contrast defect is 0.9%, again in good agreement with what the note linked in the previous elog tells us to expect...

Attachment 1: contrastDefect.pdf
contrastDefect.pdf
  12193   Thu Jun 16 18:42:12 2016 ranaUpdateCOCContrast as a function of RoC of ETMX

That sounds weird. frownIf the ETMY RoC is 60 m, why would you use 57.6 m in the simulation? According to the phase map web page, it really is 60.2 m.

  12192   Thu Jun 16 18:08:57 2016 JohannesUpdatePSLBefore the AOM installation

There was only one razor blade beam dump labeled for atmospheric use left, but that's all we need. Steve is working on restocking. I placed the modified AOM mount on the PSL table near its intended location (near the AOM where it doesn't block any beams).

Things to keep in mind:

  • The laser power needs to be turned down for the installation of the AOM. Current laser settings are: Crystal Temperature: 29.41 C, Diode Current: 2.1 A.
  • The AOM driver must not be left unterminated when turned on (which it currently is and will be).
  • The HEPA filters are currently running at ~50%. While the PSL enclosure is open for the work we'll set them to 100% and lower them after a job well done.

The setup:

The AOM has a deflection angle of about 20 mrad, which requires about 10cm of path for a separation of 2mm of the two beams. I need to survey closer and confirm, but I hope I can fit the beam dump in before the PMC (this of course also depends on the spot size). Alternatively, the PMC hopefully isn't resonant for anything remotely relevant at 80MHz offset, in which case we can also place the beam dump in its reflection path.

So this is the plan:

  • Determine coupling efficiency into PMC for reference
  • Turn down laser power
  • No signal on AOM driver modulation input
  • Mount AOM, place in beam path, and align
  • Correct alignment into PMC?
  • Secondary beam detectable? Adjust modulation input and laser power until detectable.
  • Find a place for beam dump
  • Confirm that primary beam is not clipping with PMC
  • Turn up laser power
  • Determine coupling efficiency with restored power to compare

Any thoughts? Based on the AOMs resting place I assumed that it is supposed to be installed before the PMC, but I'm actually not entirely sure where it was sitting before.

  12191   Thu Jun 16 16:11:11 2016 jamieUpdateCDSupgrade aborted for now

After poking at the new configuration more, it also started to show instability.  I couldn't figure out how to make test points or excitations available in this configuration, and adding in the full set of test point channels, and trying to do simple things like plotting channels with dtt, the frame writer (fw) would fall over, apparetnly unable to keep up with the broadcast from the dc.

I've revered everything back to the old semi-working fb configuration, and will be kicking this to the CDS group to deal with.

  12190   Thu Jun 16 15:57:46 2016 gautamUpdateCOCContrast as a function of RoC of ETMX

Summary

In a previous elog, I demonstrated that the RoC mismatch between ETMX and ETMY does not result in appreciable degradation in the mode overlap of the two arm modes. Koji suggested also checking the effect on the contrast defect. I'm attaching the results of this investigation (I've plotted the contrast, C = \frac{P\mathrm{_{max}}-P\mathrm{_{min}}}{P\mathrm{_{max}}+P\mathrm{_{min}}}  rather than the contrast defect 1-C).

Details and methodology

  • I used the same .kat file that I had made for the current configuration of the 40m, except that I set the reflectivities of the PRM and the SRM to 0. 
  • Then, I traced the Y arm cavity mode back to the node at which the laser sits in my .kat file to determine what beam coming out of the laser would be 100% matched to the Y arm (code used to do this attached)
  • I then set the beam coming out of the laser for the subsequent simulations to the value thus determined using the gauss command in finesse.
  • I then varied the RoC of ETMX (I varied the sagittal and tangential RoCs simultaneously) between 50m and 70m. As per the wiki page, the spare ETMs have an RoC between 54 and 55m, while the current ETMs have an RoC of 60.26m and 59.48m for the Y and X arms respectively (I quote the values in the "ATF" column). Simultaneously, at each value of the RoC of ETMX, I swept the microscopic position of the ETMX HR surface through 2pi radians (-180 degrees to 180 degrees) using the phi functionalilty of finesse, while monitoring the power at the AS port of this configuration using a pd in finesse. This guarantees that I sweep through all the resonances. I then calculate the contrast using the above formula. I divided the parameter space into a grid of 50 points for the RoC of ETMX and 1000 points for the microscopic position of ETMX. 
  • I fixed the RoC of ETMY as 57.6m in the simulations... Also, the maxtem option in the .kat file is set to 4 (i.e. higher order modes with indices m+n<=4 are accounted for...)

Result:

Attachment #1 shows the result of this scan (as mentioned earlier, I plot the contrast C and not the contrast defect 1-C, sorry for the wrong plot title but it takes ~30mins to run the simulation which is why I didn't want to do it agian). If the RoC of the spare ETMs is about 54m, the loss in contrast is about 0.5%. This is in good agreement with this technical note by Koji - it tells us to expect a contrast defect in the region of 0.5%-1% (depending on what parameter you use as the RoC of ETMY). 

Conclusion:

It doesn't seem that switching out the current ETM with one of the spare ETMs will result in dramatic degradation of the contrast defect...

Misc notes:

  1. Regarding the phase command in Finesse - EricQ pointed out that the default value of this is 3, which as per the manual could give unphysical results sometimes. The flags "0" or "2" are guaranteed to yield physical results always according to the manual, so it is best to set this flag appropriately for all future Finesse simulaitons. 
  2. I quickly poked around inside the cabinet near the EX table labelled "clean optics" to see if I could locate the spare ETMs. In my (non-exhaustive) search, I could not find it in any of the boxes labelled "2010 upgrade" or something to that effect. I did however find empty boxes for ETMU05 and ETMU07 which are the ETMs currently in the IFO... Does anyone know if I should look elsewhere for these?
    EDIT 17Jun2016: I have located ETMU06 and ETMU08, they are indeed in the cabinet at the X end...
  3. I'm attaching a zip file with all the code used to do this simulation. The phase flag has been appropriately set in the (only) .kat file. setLaserQparam.py was used to determine what beam parameter to assign to be perfectly matched to the Y arm. modeMatchCheck_ETM.py was used to generate the contrast as a function of the RoC of ETMX.
  4. With regards to the remaining checks to be done - I will post results of my investigations into the HOM scans as a function of the RoC of the ETMs and also the folding mirrors shortly... 
Attachment 1: contrastDefect.pdf
contrastDefect.pdf
Attachment 2: finesseCode.zip
  12189   Thu Jun 16 12:06:59 2016 ericqUpdateLSCRF Amp installed at POY11 RF output

I have installed a ZFL-500LN on the RF output of POY11. This should reduce the effect of the CM board voltage offsets by increasing the size of the error signal coming into the board. Checking with an oscilloscope at the LSC rack, the single arm PDH peak to peak voltage was something like 4mV, now it is something like 80mVyes

The setup is similar to the REFL165 situation, but with the amplifier in proximity with the PD, instead of at the end of a long cable at the LSC rack. 

The PD RF output is T'd between an 11MHz minicircuits bandpass filter and a 50 Ohm terminator (which makes sure that signals outside of the filter's passband don't get reflected back into the PD). The output of the filter is connected directly to the input of the ZFL-500LN, which is powered (temporarily) by picking off the +15V from the PD interface cable via Dsub15 breakout. (I say temporarily, as Koji is going to pick out some fancy pi-filter feedthrough which we can use to make a permanent power terminal on the PD housing.)

The max current draw of this amplifier is 60mA. Gazing at the LSC interface (D990543), I think the +15V on the DSUB cable is being passed from the eurocard crate; I don't see any 15V regulator, so maybe this is ok...

The free swinging PDH signal looked clean enough on a scope. Jamie is doing stuff with the framebuilder, so I can't look at spectra right now. However, turning the POY whitening gain down to +18dB from +45dB lets the Y arm lock on POY with all other settings nominal, which is about what we expect from the nominal +23dB gain of the amplifier.

I would see CM board offsets of ~5mV before, which was more a little more than a linewidth before this change. Now it will be 5% of that, and hopefully more manageable.

  12188   Thu Jun 16 11:25:00 2016 JohannesUpdateLSCY-Arm round-trip loss measurement with ALS

Using the ALS green beat and armlength feedback I mapped an IR resonance of the Y-Arm by stepping through a ramp of offset values.

First I optimized the IR alignment with the dither scripts while LSC kept the arm on resonance, and then transitioned the length control to ALS. The beat frequency I obtained between the Y-arm green and the PSL was about 25 MHz. Then I applied a controlled ramp signal (stepping through small offset increments applied to LSC-ALSY_OFFSET, while logging the readback from channels LSC-TRY_OUT16 and ALS-Y_FC_SERVO_INMON with an averaging time of 1s.

The plots show the acquired data with fits to  T(x)=\frac{T_0}{1+\frac{(x-x_0)^2}{\mathrm{HWHM}^2}}+\mathrm{offset} and f(x)=mx+b, respectively.

 

The fits, weighted with inverse rms uncertainty of the data points as reported by the cds system, returned HWHM = 0.6663 ± 0.0013 [offset units] and m = -0.007666 ± 0.000023 [MHz/offset unit], which gives a combined FWHM = 10,215 ± 36 Hz. The error is based purely on the fit and does not reflect uncertainties in the calibration of the phase tracker.

This yields a finesse of 388.4 ± 1.4, corresponding to a total loss (including transmissivities) of 16178 ± 58 ppm. These uncertainties include the reported accuracies of FSR and phase tracker calibration from elog 9804 and elog 11761.

The resulting loss is a little lower than that of elog 11712, which was done before the phase tracker re-calibration. Need to check for consistency.

Attachment 1: T_vs_steps.pdf
T_vs_steps.pdf
Attachment 2: f_vs_steps.pdf
f_vs_steps.pdf
  12187   Thu Jun 16 11:10:17 2016 SteveUpdateSUSspare ETMX optics preparation to be hanged

D Location

Number on

Drawing

 

Component

Name

 

 Baked Clean

 Pieces Needed

 

Pieces

In Stock

(on tower )

Notes
         
31,20,19,13 viton tips   not cut yet baked material in stock
30 6-32x0.75" stops 4 (4)  
29 steel music wire 0.017" not baked on roll  needs good wipes
28  1/4 "washer 4 (4)  
27 lock washer 4 50 install
26 Ag plated 1/4-20x1.25 4 (4)  
25 Ag plated1/4-25x.75 & not plated 20 (20)  
24 SS 4-40x.5 2 (2)  
23 SS 4-40x.38 4 (4)  
22 spring plunger 4+2 (4+2)  
         
         
18 magnets, 1.9mm od, length 3.2 mm 5 ~30 not coated, rusty

buy Ni coated ones for future use from www.electroenergy.com

 

17 guide rod, 0.635 mm od, 3.3 mm 3 6 Al
16 wire standoff, 1 mm od, 4.8 mm 2 2 Al and ruby (ruby groove not centered)
15 short OSEMs 5 6  
14 spare ETMX in 40m wiki 1 1  confirmed in cabinet
         
12 dumbbell standoff 5 6  
11 Al stiffening plate 1 (1)  
10 wire clamp B in sus block 2 (2)  
9 wire clamp A 1 (1)  
8-7 lower clamp for lifting optics 2 (2)  
6 upper clamp to hold down optic 1 (1)  
5-4 left-right side of tower  1 ea (1ea)  
3 tower base 1 (1)  
2 sus block 1 (1)  
1 lower and upper OSEM holders 1ea (1ea)  
         
48 sandind fixture for magnet&dumbbell      
45 magnet-dumbbell assemblly fixture 1 1  
43 guide-wirestandoff gluing fixture 1 1

Ok for larger RUBY,

unit is not in perfect condtition but usefull

 

         
  First contact   3-15-2013 purchase

Pick up  FC from Gary with purchasing date 7-7-2015

or later

  FCPEEK peeler ring disk for TM cleaning 10 front,10 back side  have sheets only 32&19mm ID punches ordered
  GordonBrush custom for LIGO optical cleaning ~5 1 (3/8wide nylonSS) more from Calum available
  EP30-2 epoxy have  have expiration date 9-24-2016

NOT finished, last edited 7-7

Attachment 1: SOSsus_Page_1.png
SOSsus_Page_1.png
Attachment 2: SOSsus_Page_2.png
SOSsus_Page_2.png
Attachment 3: ETMXspareTowerFace.jpg
ETMXspareTowerFace.jpg
Attachment 4: ETMXspareTowerBack.jpg
ETMXspareTowerBack.jpg
Attachment 5: magnets-dumbbells.jpg
magnets-dumbbells.jpg
ELOG V3.1.3-